# SpiralShard: Highly Concurrent and Secure Blockchain Sharding via Linked Cross-shard Endorsement

### v1

@@ IV-B Efficient Cross-shard Communication for LCE
To verify the validity of a prepared block before endorsing, a straw-man design is to store the shard state and verify each raw transaction inside that block, leading to massive overhead. However, we hope to complete the verification before endorsement with little overhead. To achieve this goal, we fine-tune the system parameters according to the mathematical expressions in Section V-A to ensure the fraction of malicious nodes in each shard is less than $2/3$, the quorum size of BFT-typed consensus, with a high probability. Each prepared block is valid since it has been verified and voted for by at least one honest node in its shard. Moreover, we require each node to be responsible for the cross-shard communication of the blocks it votes for. Hence, at least one honest node will broadcast the prepared block to the corresponding shard for subsequent endorsement and finalization. With the nodes’ identity recorded in the beacon chain, each node from another shard only needs to verify whether the received block is prepared by verifying the signatures in the block header. Therefore, we reuse the block header (instead of the whole block) with two additional fields for endorsement and finalization. 1) The latest finalized hash indicates the hash of the latest finalized block in its shard. 2) The endorsement list consists of hashes of up to $M-1$ blocks going back to the original shard’s block, indicating the recent history of the linked endorsements. The endorsement list is updated across each step of the linked endorsements. For example, when a shard proposes block $B$ to endorse the latest block $A$ from the preceding shard in the endorsement group, block $B$’s endorsement list is copied from that in block $A$ with an additional hash of block $A$.
@@ IV-C Pipelining Mechanism
IV-D Other Components
Beyond our main designs, we provide brief remarks about implementing the other components, including bootstrapping, shard reconfiguration, cross-shard transfer transactions, and cross-shard smart contracts.
First, before the first epoch, there is a one-time trusted bootstrapping process, a reasonable assumption widely adopted in research and industrial areas [5, 29]. In practice, bootstrapping can be carried out in blockchain projects in a centralized [5] or decentralized way [46].
Second, since the adversary is slowly-adapted (i.e., it can infect some honest nodes after a while), we periodically reshuffle shards to prevent aggregation of malicious nodes. Between two epochs, SpiralShard has a shard reconfiguration process (also existing in most existing works [46, 21, 47, 5]) to determine members of each shard in the next epoch. For security, the nodes are randomly assigned to different shards. So we leverage a combination of verifiable random function (VRF) [37] and verifiable delay function (VDF) [9] to generate verifiable, unbiased, and unpredictable distributed randomness for this process, similar to that of other sharding systems [29, 46, 21]. To prevent Sybil attack, SpiralShard applies Proof of Stake (PoS) as the access mechanism for nodes who want to join the system. Similarly, blockchain systems can use PoW as an access scheme [46, 21].
Third, similar to many existing works [21, 29, 5], there is a beacon chain in SpiralShard that records the nodes’ identity information and assists in shard reconfiguration. Note that the beacon chain maintains the identity of participants for each epoch, epoch randomness, and the starting block height of each epoch, which assists the progress of LCE across epochs and will be discussed later in Section V-C. Thus, the identity of a node for each epoch is public.
Fourth, the cross-shard transfer transactions are processed via the traditional relay-based mechanism [5, 45]. In SpiralShard, each cross-shard transfer transaction involves two accounts in different shards. Specifically, source shards are responsible for initially packaging cross-shard transactions to complete deduction operations. Packaged transactions and proof of deduction operations’ finalization are routed to the destination shards via Kademlia routing protocol [46, 5]. Upon receiving the cross-shard transactions and proof from the source shards, the destination shard will package the transactions to complete the deposit operations. To prove the finalization of a cross-shard transaction in the source shard, nodes must send two items. First, they need to send the Merkle proof [14], which validates the path from the transaction to the transaction Merkle tree’s root within the block header’s metadata, similar to existing works like [21, 26, 45, 5]. Second, they must provide a subsequent block header from the source shard, proving the transactions’ block has been finalized. Accordingly, the destination shard will verify whether the cross-shard transaction is finalized in the source shard before completing the deposit operation in the receiver’s account, similar to existing works [21, 26, 45, 5]. Note that our system does not diminish the foundational usability of the cross-shard transaction processing scheme, even if two accounts are in different endorsement groups. The rationale is that shard identifiers of nodes are visible to the entire network, so even if the shards are not in the same endorsement group, they can mutually verify the signature in block headers.
Fifth, our system is inherently capable of supporting smart contracts, primarily because we ensure the safe finalization of blocks. To enable cross-shard contract transactions, we can make slight modifications inspired by the Two-Phase Commit (2PC) protocol from the SGX-Sharding [17]. Specifically, a shard packages a BeginTX into a block when a client requests to call its smart contract. Once this block is finalized, the shard sends the BeginTX with finality proof (Merkle tree path and block header) to relevant shards to lock states. Upon verifying the proof, corresponding shards package PrepareOK messages to lock states (or PrepareNotOK messages in cases where locking isn’t feasible). The finalized PrepareOK/PrepareNotOK messages are sent back to the primary shard with finality proof. This primary shard will package a CommitTX to execute the original smart contract if all the required states are locked (or AbortTX messages otherwise). Once the CommitTX/AbortTX is finalized, it is broadcast back to the corresponding shards with the proof of finality to execute the commit or abort operation. It is worth noting that every message (BeginTX, PrepareOK, PrepareNotOK, CommitTX, and AbortTX) is transmitted with an accompanying finality proof to preserve atomicity.
V Analysis
V-A Epoch Security
@@
In this section, we calculate the upper bound of the failure probability within each epoch. In SpiralShard, we leverage cumulative hypergeometric distribution function to calculate probabilities, similar to previous works [46, 21, 31, 47, 29]. Besides, previous works typically assume the failure probability of each shard is independent while adopting union bound [^2] The union bound delivers an upper limit or a ’worst-case scenario’ estimate for the occurrence of any event within a given set. to calculate the upper bound of the system’s failure probability [46, 21, 31, 47, 29]. We adopt the same assumption and mathematical tools. In SpiralShard, an endorsement group will fail in three cases: Case 1: It contains no honest shard, i.e., all shards in this endorsement group have at least $1/3$ fraction of malicious nodes. Case 2: It contains honest shards, but at least one shard is corrupted by at least $2/3$ fraction of malicious nodes. Case 3: It contains honest shards, but at least one shard is corrupted by at least $1/3$ fraction of Byzantine nodes.
@@
$Pr[\text{Failure}_{i}^{(\text{Case 1})}]\leq Pr[X_{i}\geq M/3]=\sum_{x=\lfloor M/3\rfloor}^{M}Pr[X_{i}=x].$ (2)
@@
Ensuring Negligible Failure Probability. SpiralShard must ensure a negligible failure probability within each epoch to maintain security, similar to most blockchain sharding works [29, 46, 47, 21, 31]. Based on Equations 14, and 8, we need to adjust the shard size $S$ and endorsement group size $G$ to make sure there is a small $\varepsilon$ existing so that:
@@
The specific values of the probability can be found in Table I.
V-B Protocol Security Analysis
Based on negligible epoch failure probability, we now analyze the safety and liveness of our LCE protocol. Specifically, safety indicates that honest nodes agree on the sequence of the same and valid finalized blocks in each shard chain. Liveness indicates each shard continuously has blocks being finalized. Thus, the system progresses, similar to other permissionless sharding systems [46, 29]. We first analyze the safety of the LCE protocol.
@@ V-B Protocol Security Analysis
The LCE protocol provides safety if each corrupted shard has less than $2/3$ malicious nodes within the same epoch and each endorsement group has less than $1/3$ of malicious nodes within the same epoch.
@@
Without loss of generality, we demonstrate the safety of a specific finalized block $B_{a}$. Since block $B_{a}$ is finalized, it has been verified by at least one honest node and has received endorsements from all shards in its endorsement group within the epoch. Assuming another block, $B_{b}$, which conflicts with $B_{a}$, is finalized in the same epoch. This assumption is based on two distinct scenarios. Firstly, all shards in the endorsement group simultaneously endorse these two blocks. In this case, the proportion of malicious nodes in the endorsement group must be at least $1/3$, as each shard within the group is corrupted by at least $1/3$ fraction of malicious nodes. However, it contradicts the assumption that the fraction of malicious nodes within each endorsement group is less than $1/3$. Secondly, a corrupted shard manipulates the other shards’ endorsement decisions by setting an invalid endorsement list in its block header. This requires the number of malicious nodes within this corrupted shard to be larger than the quorum size of BFT-typed consensus. However, it contradicts the assumption that the fraction of malicious nodes within each shard is less than $2/3$. In both scenarios, the assumptions of the theorem are violated. Thus, if the conditions of the theorem are met, the safety is preserved without losing generality. ∎
We first prove that the intra-shard consensus is continuously reached in Lemma 1, similar to the proof of liveness in Flexible BFT [36]. Based on this, we then prove that the LCE protocol can preserve liveness.
Lemma 1.
The intra-shard BFT consensus is continuously reached if there are less than $1/3$ fraction of Byzantine nodes plus no more than $1/3$ fraction of a-b-c nodes for each shard and at least one honest shard within each endorsement group.
@@
In accordance with the definition of a-b-c nodes, elaborated in Section III-C and supported by the existing work [36], these nodes preserve liveness. Indeed, as demonstrated in Theorem 1, safety is provided if each endorsement group has at least one honest shard and no shard corrupted by at least $2/3$ of malicious nodes. Consequently, when proving Lemma 1 and once safety is established, the a-b-c nodes would not prevent intra-shard consensus being reached aligning with the principles of a-b-c nodes in the existing work [36]. In addition, BFT-typed consensus protocols typically have a view-change mechanism that relies on the participation of $2/3$ nodes to change leaders when no consensus is reached (e.g., PBFT [12] and FBFT [5], the variant of PBFT). In this case, if a Byzantine leader prevents liveness (e.g., it remains silent or proposes invalid blocks that cannot pass intra-shard consensus), at least $2/3$ fraction of non-Byzantine nodes would participate in the view change to ensure a new round of consensus can be reached. Ultimately, a non-Byzantine node will assume leadership and advance intra-shard consensus. ∎
Theorem 2.
The LCE protocol provides liveness.
Proof.
According to Lemma 1, each shard reaches consensus continuously. According to Theorem 1, all the honest nodes would agree on the same finalized blocks without conflicts. Moreover, the endorsement is sequentially made to prevent different honest shards from endorsing blocks in conflicting fork branches. In this case, each shard only needs to endorse one previous shard directly and can implicitly endorse the endorsement decisions of the preceding shard. So honest shards will make non-conflicting endorsements. Therefore, the system will not be blocked indefinitely due to conflicts between honest shards.
However, a corner case is that a corrupted shard’s Byzantine leader refuses to endorse any endorsable block at its proposed block to halt the LCE. In this case, similar to Lemma 1, non-Byzantine (a-b-c and honest) nodes will participate in view change until the newly proposed block does the endorsement correctly. In the worst-case scenario outlined in BFT2F [30], multiple view changes may succeed simultaneously if there are less than $2/3$ of malicious nodes, leading to forking. Fortunately, LCE will eliminate the forks by providing safe finalization. ∎
V-C Cross-epoch Scenario
During epoch transitions, each shard’s members are changed, resulting in a block produced near the end of an epoch that $M$ corrupted shards may endorse. In this case, a finalized block may be unsafe, i.e., only corrupted shards endorse it at the end of one epoch, and other shards in this endorsement group are corrupted at the next epoch. To tackle this vulnerability, the finalization of the block requires the accumulation of endorsements within the same epoch. This additional verification can be done since each node can access the shard’s participants and the block height for the next epoch from the beacon chain. Additionally, if a block is not sufficiently endorsed before the epoch ends, it will remain prepared and can still be endorsed to preserve liveness. However, it must be endorsed from scratch in the next epoch before being finalized.
V-D Pipelining Mechanism Analysis
The pipelining mechanism is designed to maintain throughput with the existence of some slow shards. It requires each shard to prepare blocks optimistically and endorse the latest block of the preceding shard in the endorsement group order. In this case, for a shard reaching consensus at a higher speed, not all blocks can be endorsed by the succeeding shard. Fortunately, once an optimistically prepared block is sufficiently endorsed, all the parent blocks in this shard chain are safely finalized. The rationale is that blocks finalized by the LCE protocol will not conflict, thus forming a single chain without any forks. Besides, we suggest developers reduce the size of the endorsement group under negligible epoch failure probability, as mentioned in Section VI-B, to reduce the impact of slow shards on latency.
V-E Corner Case and Limitations
Apart from typical attacks (i.e., transaction manipulation, double-spending, and silent attack), DDoS attacks may be launched by Byzantine nodes or nodes outside the blockchain network to break liveness. Since our network model requires every message to be signed by nodes within the network, unsigned messages will be treated as spam messages and discarded by the system. Besides, Byzantine leaders of corrupted shards may propose too many forking blocks to consume other nodes’ network bandwidth, compromising the liveness. Fortunately, nodes can generate a witness of such an attack and propose leader rotation.
Furthermore, the potential exists for malicious leaders to initiate censorship attacks, and the current system cannot resist such attacks. Besides, a malicious leader may determine the order of the transaction in a block for extra revenues (a.k.a. miner extracted value, MEV [16]). Fortunately, leaderless BFT-typed consensus is capable of dealing with that. For instance, Red Belly [13] involves merging micro-blocks proposed by multiple nodes into a unified block, thus ensuring the integrity of transactions. Additionally, this consensus protocol offers defense against eclipse attacks targeting specific leaders. Notably, related research is orthogonal to our work and will be considered adapted as the pluggable intra-shard consensus mechanism.
Finally, the cross-chain execution of smart contracts is naturally supported by SpiralShard. The rationale is that SpiralShard provides secure and highly concurrent confirmations for blocks. However, to ensure atomicity, it is required that transitional information (i.e., BeginTX, PrepareOK/PrepareNotOK, CommitTX/AbortTX) must be finalized via LCE protocol before being transmitted across shards (as mentioned in Section IV-D). This imposes a limitation that multiple rounds of finalization delays should be considered. However, given that our system’s primary objective is to enhance concurrency, we primarily focus on transfer transactions. The efficiency optimization for smart contract execution remains a consideration for future work.
VI Evaluation
VI-A Implementation and Experimental Setup
We implement a prototype of SpiralShard in Golang based on Harmony [5], a well-known permissionless blockchain sharding project once with top 50 market cap in cryptocurrency. We choose FBFT [5] consensus protocol proposed by Harmony as our intra-shard consensus protocol. Such a choice is made for 1) simplifying the implementation process and 2) providing a fair evaluation to show the performance gains from the main ideas of SpiralShard.
We deploy a network with up to $4,200$ nodes on $21$ Amazon EC2 instances, each having a 96-core CPU and up to 20 Gbps network bandwidth. We limit the network by using Docker container and Linux Traffic Control (TC). Each node has 50 Mbps individual network bandwidth and 100 ms latency for each message. We set the transaction size as $512$ bytes and let each block contain up to $4,096$ transactions. Moreover, we generate transactions based on historical data of Ethereum [48], where cross-shard transactions’ proportion increases with the number of shards. The system’s total fraction of malicious nodes is set to $1/4$, including $1/8$ fraction of Byzantine nodes and $1/8$ fraction of a-b-c nodes.
VI-B Parameter Settings
In SpiralShard, the endorsement group and shard sizes are the primary system parameters that are fine-tuned while running. In our implementation, we leverage an authorized account (controlled by developers) that can publicly set one of these parameters in the beacon chain. Note that the setting can be decentralized by using a Decentralized Autonomous Organization (DAO). When one of the two parameters is determined, each node can individually adjust the other parameter from small to large until a negligible failure probability is obtained to get the exact result. We require the calculated upper bound to be less than $2^{-17}\approx 7.6\times 10^{-6}$, meaning that the system fails once in 359 years, the same as the existing systems [21, 31].
In the baseline, we set the shard size under the negligible failure probability of the system according to the classical hypergeometric distribution function (e.g., the equations in [46]), as the baseline does not have our LCE design.

### v2

@@ V-B Efficient Cross-shard Communication for LCE
To verify the validity of a prepared block before endorsing, the aforementioned basic design requires to store the shard state and verify each raw transaction inside that block, leading to massive overhead. However, we hope to complete the verification before endorsement with little overhead.
To achieve this goal, our core idea is to let each shard ensure the validity of individual transactions through intra-shard consensus, while employing the LCE protocol to resolve potential forks at the shard level using only block headers. In this way, the resource-intensive task of verifying specific transactions based on shard state remains entirely within each shard. Meanwhile, the LCE protocol leverages lightweight information contained in the block headers (e.g., consensus signatures) to efficiently detect and resolve forks introduced by malicious shards.
Specifically, we fine-tune the system parameters according to the mathematical expressions in Section VI-A to ensure the fraction of malicious nodes in each shard is <2/3<2/3, the quorum size of BFT-typed consensus, except with negligible failure probability. In this scenario, the adversary cannot control a sufficient number of malicious nodes to approve blocks containing invalid transactions through intra-shard consensus. Specifically, when the fraction of malicious nodes is below the required quorum threshold for intra-shard consensus (i.e., 2/3 of the shard size), any attempt to reach consensus on a block containing invalid transactions—thus achieving the quorum threshold (2/3)—necessitates obtaining at least one honest node’s signature. However, honest nodes verify transaction validity rigorously and consequently will never endorse such invalid transactions. Therefore, under this condition, all prepared blocks that successfully pass intra-shard consensus must contain only valid transactions.
However, even though malicious nodes constituting less than 2/3 of the shard are unable to tamper with transactions (as above-mentioned), they can still induce forks within the shard to undermine security. Fortunately, the detection of such forks can be efficiently achieved through block headers. Specifically, with the nodes’ identity recorded in the beacon chain, each node from another shard only needs to verify whether the received block is prepared by verifying the signatures in the block header. Moreover, we require each node to be responsible for the cross-shard communication of the blocks it votes for. Hence, at least one honest node will broadcast the prepared block to the corresponding shard for subsequent endorsement and finalization. As a result, we can reuse the block header (instead of the whole block) with two additional fields for endorsement and finalization. 1) The latest finalized hash indicates the hash of the latest finalized block in its shard. 2) The endorsement list consists of hashes of up to M1M-1 blocks going back to the original shard’s block, indicating the recent history of the linked endorsements. The endorsement list is updated across each step of the linked endorsements. For example, when a shard proposes block $B$ to endorse the latest block $A$ from the preceding shard in the endorsement group, block $B$’s endorsement list is copied from that in block $A$ with an additional hash of block $A$.
@@ V-C Pipelining Mechanism
VI Analysis and Discussion
VI-A Epoch Failure Probability Analysis
@@
In this section, we calculate the upper bound of the failure probability within each epoch. In SpiralShard, we leverage cumulative hypergeometric distribution function to calculate probabilities, similar to previous works [rapidchain, pyramid, li2022jenga, cycledger, omniledger]. Besides, previous works typically assume the failure probability of each shard is independent while adopting union bound [^2] The union bound delivers an upper limit or a ’worst-case scenario’ estimate for the occurrence of any event within a given set. to calculate the upper bound of the system’s failure probability [rapidchain, pyramid, li2022jenga, cycledger, omniledger]. We adopt the same assumption and mathematical tools. In SpiralShard, an endorsement group will fail in three cases: Case 1: It contains no honest shard, i.e., all shards in this endorsement group have at least $1/3$ fraction of malicious nodes. Case 2: It contains honest shards, but at least one shard is corrupted by at least $2/3$ fraction of malicious nodes. Case 3: It contains honest shards, but at least one shard is corrupted by at least $1/3$ fraction of Byzantine nodes.
@@
[…]
@@
Ensuring Negligible Failure Probability. SpiralShard must ensure a negligible failure probability within each epoch to maintain security, similar to most blockchain sharding works [omniledger, rapidchain, cycledger, pyramid, li2022jenga]. Based on Equations 14, and 8, we need to adjust the shard size $S$ and endorsement group size $G$ to make sure there is a small $\varepsilon$ existing so that:
@@
The specific values of the probability can be found in Table II.
VI-B Protocol Security Analysis
Based on negligible epoch failure probability, we now analyze the safety and liveness of our LCE protocol.
@@ VI-B Protocol Security Analysis
The LCE protocol provides safety if each corrupted shard has less than $2/3$ malicious nodes (i.e., less than $1/3$ fraction of Byzantine nodes plus no more than $1/3$ fraction of a-b-c nodes) and at least one honest shard exists in each endorsement group (i.e., each endorsement group has less than $1/3$ of malicious nodes) within the same epoch.
@@
We use proof by contradiction to demonstrate this theorem. Without loss of generality, we demonstrate the safety of a specific finalized block $B_{a}$. Since block $B_{a}$ is finalized, it has been verified by at least one honest shard and has received endorsements from all shards in its endorsement group within the epoch. Assuming another block, $B_{b}$, which conflicts with $B_{a}$, is finalized in the same epoch. This assumption is based on two distinct scenarios. First, all shards in the endorsement group simultaneously endorse these two blocks via intra-shard consensus. This means, in this case, there is no honest shard in the endorsement group (i.e., the proportion of malicious nodes in the endorsement group is at least $1/3$). However, this contradicts the theorem that at least one honest shard exists in each endorsement group. Second, a corrupted shard manipulates the other shards’ endorsement decisions by setting an invalid endorsement list in its block header during intra-shard consensus. This requires the number of malicious nodes within this corrupted shard to be larger than the quorum size of BFT-typed consensus. However, it contradicts the theorem that the fraction of malicious nodes within each shard is less than $2/3$. In both scenarios, the theorem is violated. Therefore, if the conditions of the theorem are met, the safety is preserved without losing generality.
Notably, the presence of malicious nodes, which exceed 1/3 but remain below 2/3 in a shard, can potentially introduce nondeterminism during the view change process within corrupted shards. For instance, due to the presence of a malicious leader or a large number of malicious nodes, these nodes may deliberately manipulate information, causing some honest nodes to believe that the view change has been completed while others do not perceive a need for it. This may lead to inconsistent states among honest nodes and the persistent formation of forks. In SpiralShard, we tolerate such nondeterminism within corrupted shards, as forks resulting from nondeterminism are ultimately resolved by the LCE protocol, which selects a single branch to finalize (Section V-A). Once a fork in a corrupted shard is finalized through the LCE protocol, honest nodes within the shard that had previously chosen a different branch (due to nondeterminism) will roll back their state to the latest finalized state determined by the LCE protocol. They will then proceed with state updates based on the branch selected by the current LCE protocol. Overall, any nondeterminism arising from intra-shard consensus or view change processes within corrupted shards can eventually be resolved by the LCE protocol, which finalizes only one of the forks.
Theorem 2.
The LCE protocol provides liveness if each corrupted shard has less than $1/3$ fraction of Byzantine nodes plus no more than $1/3$ fraction of a-b-c nodes and at least one honest shard exists in each endorsement group within the same epoch.
@@
We first prove that the intra-shard BFT-typed consensus can maintain liveness, and then, based on this, prove that the LCE protocol can maintain liveness.
First, for the intra-shard consensus. In accordance with the definition of a-b-c nodes, elaborated in Section III-B and supported by the existing work [flexible], these nodes preserve liveness. Indeed, as demonstrated in Theorem 1, safety is provided. Consequently, once safety is established, the a-b-c nodes would be treated as honest and would not prevent intra-shard consensus (also view change) process from being proceeded, aligning with the principles of a-b-c nodes in the existing work [flexible]. Since no more than 1/3 of the nodes in each shard are a-b-c nodes, and these nodes do not attack liveness, the proportion of nodes attacking liveness in each shard remains below 1/3, which is within the fault tolerance threshold of BFT-typed consensus. Therefore, within each shard, both the consensus process and the view change process can maintain liveness.
Second, for the LCE protocol. As mentioned above, the consensus and view change processes within each shard maintain liveness. This means that each shard can continuously endorse other shards through its consensus process. Furthermore, based on Theorem 1 and Section V-A, an honest shard will always select a single fork for endorsement through intra-shard consensus. Given these two points, the LCE protocol requires that shards sequentially endorse the preceding shard in each endorsement group. In this case, each shard only needs to endorse one previous shard directly and can implicitly endorse the endorsement decisions of the preceding shard. When a shard endorses, it is aware of which fork its preceding shard has endorsed, allowing honest shards to make non-conflicting endorsements. Therefore, in the LCE protocol, the sequential endorsements between shards prevent situations where different honest shards endorse different forks, leading to insufficient endorsements for any particular fork. Consequently, through the LCE protocol, honest shards within each endorsement group sequentially endorse the same fork. This ensures that, within the same endorsement group, there will always be one fork that collects endorsements from all shards, thereby guaranteeing the liveness of the LCE protocol.
VI-C Corner Cases Discussion
Apart from typical attacks (i.e., transaction manipulation, double-spending, silent attack), DDoS attacks may be launched by Byzantine nodes or attackers outside the blockchain network to break liveness. For Byzantine nodes within the system, for example, Byzantine leaders may propose too many forking blocks to consume other nodes’ network bandwidth, compromising the liveness. Fortunately, many robust BFT protocols today (such as Tendermint [buchman2016tendermint] or Harmony’s FBFT [Harmony]) typically incorporate multiple mechanisms to resist or mitigate such attacks. For instance, nodes can apply bandwidth or rate limits to incoming blocks, or use ’gossip by demand,’ allowing nodes to request blocks from the leader only when needed, rather than having the leader proactively broadcast all blocks. Additionally, each node can quickly detect whether the leader has sent multiple blocks through signature verification. Upon detecting equivocation, nodes can discard all subsequent messages from that leader. Regarding external attackers, such as those attempting to launch attacks by submitting numerous spam transactions or sending a large number of invalid data packets to blockchain nodes, there are several classic defense measures. Examples include employing transaction fees to increase attack costs, applying traffic control to RPC nodes, or using DDoS protection tools (e.g., Cloudflare, AWS Shield) to protect the external IP addresses of nodes.
Censorship attacks represent another possible threat, where a malicious leader selectively chooses not to include specific transactions in the blocks [mev]. Such attacks can be mitigated through various additional measures. For instance, privacy-enhancing technologies [javed2021petchain] can be used to conceal the content of transactions, making them less susceptible to censorship. Alternatively, leaderless BFT-type consensus protocols can be employed. For instance, Red Belly [redBelly] involves merging micro-blocks proposed by multiple nodes into a unified block, thus ensuring the integrity of transactions. Additionally, this consensus protocol offers defense against eclipse attacks targeting specific leaders.
In summary, most of the aforementioned attacks, along with other potential threats, can typically be mitigated through well-established standard defense mechanisms. As these mechanisms are specifically designed to counter their respective attack vectors, they help ensure that system performance remains largely unaffected. Importantly, these lines of research are orthogonal to the core design of SpiralShard. Nonetheless, SpiralShard can be readily integrated with such existing solutions to enhance its resilience against atypical attacks while maintaining system performance.
VI-D Cross-shard Smart Contract
For cross-shard smart contract processing, we make slight modifications inspired by the Two-Phase Commit (2PC) protocol [sgxSharding]. Specifically, a shard packages a BeginTX into a block when a client requests to call its smart contract. Once this block is finalized, the shard sends the BeginTX with finality proof (Merkle tree path and block header) to relevant shards to lock states. Upon verifying the proof, corresponding shards package PrepareOK messages to lock states (or PrepareNotOK messages in cases where locking isn’t feasible). The finalized PrepareOK/PrepareNotOK messages are sent back to the primary shard with finality proof. This primary shard will package a CommitTX to execute the original smart contract if all the required states are locked (or AbortTX messages otherwise). Once the CommitTX/AbortTX is finalized, it is broadcast back to the corresponding shards with the proof of finality to execute the commit or abort operation. It is worth noting that every message (BeginTX, PrepareOK, PrepareNotOK, CommitTX, and AbortTX) is transmitted with an accompanying finality proof to preserve atomicity.
One limitation of SpiralShard is that it may increase transaction confirmation latency, as (cross-shard) transactions need to wait for finalization by the LCE protocol before they can be confirmed (Section IV-C). This is especially true for cross-shard smart contract transactions, which may involve multiple rounds of finalization delays. However, as the main goal of this work is to improve concurrency, reducing transaction latency is left as a consideration for future research.
VII Evaluation
VII-A Implementation and Experimental Setup
We implement a prototype of SpiralShard in Golang based on Harmony [Harmony], a well-known permissionless blockchain sharding project once with top 50 market cap in cryptocurrency. For a balanced evaluation, we adopt Harmony as the primary baseline system (hereafter Baseline) and benchmark it against SpiralShard. Using Harmony isolates the impact of our design decisions and highlights the performance benefits attributable to SpiralShard’s core mechanisms. In fact, SpiralShard’s mechanisms can be seamlessly incorporated into many existing sharding frameworks [elastico, omniledger, rapidchain, sgxSharding, pyramid, cycledger]. In addition, we implement a prototype of GearBox [gearbox] (CCS ’22)–a representative state-of-the-art sharding protocol that explicitly tolerates corrupted shards–to provide a contemporary point of comparison in our performance study.
We deploy a network with up to $4,200$ nodes on $21$ Amazon EC2 instances, each having a 96-core CPU and up to 20 Gbps network bandwidth. Each node has 50 Mbps individual network bandwidth and 100 ms latency for each message. We set the transaction size as $512$ bytes and let each block contain up to $4,096$ transactions. The system’s total fraction of malicious nodes is set to 1/4, including 1/8 Byzantine nodes and 1/8 a-b-c nodes. In most of the following experiments, by default, these malicious nodes will try to generate a fork constantly. Specifically, a malicious leader can attempt to fork the shard’s chain whenever it proposes a block. As a result, with a 1/4 malicious node ratio, this yields a fork attempt roughly 25% of the time. Moreover, forking is not the only attack scenario evaluated. For example, Section VII-H presents a silent attack scenario, and Section VII-I evaluates performance under adversarial network latency (attackers trying to delay the network).
In our evaluation, the transactions are generated based on the historical data of Ethereum [xblock]. To be more realistic, the proportion of cross-shard transactions in the experiments increases with the number of shards. This means that, given the same network scale, SpiralShard can configure a greater number of smaller shards, resulting in a higher proportion of cross-shard transactions compared to the baseline. We choose to conduct experiments under this realistic cross-shard transaction ratio (rather than controlling the cross-shard transaction ratio to be identical between SpiralShard and the baseline) mainly to facilitate a more fair and authentic performance comparison.
VII-B Parameter Settings
The parameter settings are shown in Table II. In SpiralShard, the endorsement group and shard sizes are the primary system parameters that are configured by the developers. We require the calculated upper bound to be less than $2^{-17}\approx 7.6\times 10^{-6}$, meaning that the system fails once in 359 years, the same as the existing systems [pyramid, li2022jenga]. In the baseline (Harmony), we set the shard size under the negligible failure probability of the system according to the classical hypergeometric distribution function (e.g., the equations in [rapidchain]), as the baseline does not have our LCE design. In GearBox, we do not provide a failure probability since the framework itself lacks a methodology for calculating system-wide failure rates. Instead, we strictly follow the shard-scaling approach outlined in the GearBox paper by gradually increasing the shard size from small to large. After four such increases (as specified in their paper [gearbox]), we compute the number and size of shards that can be successfully configured in their system. Moreover, to ensure evaluation fairness, we configure each node in GearBox to participate in one shard, the same as we do.